OpenStudy (anonymous):

The original price of a television decreases by 20 percent. By what percent must the price increase to reach its original value?

OpenStudy (anonymous):

25%

OpenStudy (anonymous):

How come?

OpenStudy (anonymous):

its 20 percent

OpenStudy (anonymous):

Prove it.

OpenStudy (anonymous):

ok so what is 25% of 80?

OpenStudy (anonymous):

20,right?

OpenStudy (anonymous):

let x be price of television decrease in price=20/100*x=x/5 current price of television =x-x/5=4x/5 for this should be equal to x say k should be added implies 4x/5+k=x then k=x/5 i.e, 20%

OpenStudy (anonymous):

my mistake its 25 percent iwil prove

OpenStudy (anonymous):

Let the original price be 100 after 20% decrease it becomes 80 now to again make it 100(the original price) we need to add 20 to it thus the percentage increase will be given as (20/80)*100 hence it gives 25% so the price should be increased by 25%

OpenStudy (anonymous):

i was right

OpenStudy (anonymous):

yup i did a mistake in my calc

OpenStudy (anonymous):

sry its 25 error in typing

OpenStudy (anonymous):

Man, I don't like it when people use x/x as fraction, why not use the equation button. T_T It takes me longer to compute.

OpenStudy (anonymous):

hey shast r u happy u got the anser rite

OpenStudy (anonymous):

Lol. I don't care if the answer is right. I care on how to solve it.

OpenStudy (anonymous):

kk,i gave u the steps tooo

OpenStudy (anonymous):

understood the procedure

OpenStudy (anonymous):

I like surd steps better become its easier for me to understand with those sign.

OpenStudy (anonymous):

Though both of you got the same idea.

OpenStudy (anonymous):

thnxx shast

OpenStudy (anonymous):

No, thank you... and of course gil.

OpenStudy (anonymous):

thnxx shasat